OpenStudy (anonymous):

The blades of a windmill start from rest and rotate with an angular acceleration of 18.0 rad/s2. At any point on a blade, how much time passes before the magnitude of the tangential acceleration equals the magnitude of the centripetal acceleration? I am not sure how to work this.

OpenStudy (jamesj):

First, what are expressions for the tangential and centripetal accelerations?

OpenStudy (anonymous):

centripetal acceleration=v^2/r \[tangential= r*\alpha\]

OpenStudy (jamesj):

write the first one in terms of angular velocity terms

OpenStudy (jamesj):

\[ v = \omega r \]

OpenStudy (anonymous):

thanks, how do I incorporate time into this?

OpenStudy (jamesj):

Well, what is angular velocity as a function of time? \[ \omega(t) = ...what? \]

OpenStudy (anonymous):

oh, ok- I've got it from here, thanks. I had some brain fart on that part.

OpenStudy (jamesj):

Ok. For the record, if T(t) is the tangential acceleration, C(t) the centripetal, then \[ T(t) = r\alpha^2 t^2, \ \ \ C(t) = r\alpha \] which are equal when \( t = 1/\sqrt{\alpha} \).

OpenStudy (anonymous):

hey how did you find the function of T and C WITH RESPECT TO TIME?